### v1

# Some one-dimensional elliptic problems with constraints

Abstract.
Given $m\in\mathbb{N}\setminus\{0\}$ and $\rho>0$, we find solutions $(\lambda,u)$ to the problem
$\begin{cases}\bigl{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}\bigr{)}^{m}u+\lambda G^{\prime}(u)=F^{\prime}(u)\\ \int_{\mathbb{R}}K(u)\,\mathrm{d}x=\rho\end{cases}$
in the following cases: $m=1$ or $2G(s)=K(s)=s^{2}$. In the former, we follow a bifurcation argument; in the latter, we use variational methods.
Key words and phrases:
Poly-harmonic Schrödinger equations, homoclinic solutions, normalised solutions, least-energy solutions, bifurcation theory, variational methods
2020 Mathematics Subject Classification:
34L40, 35B32, 35J15, 35J35, 35J91
1. Introduction and statement of the results
We study the problem
(1.1) $\begin{cases}\displaystyle\Bigl{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}\Bigr{)}^{m}u+\lambda G^{\prime}(u)=F^{\prime}(u)\\ \displaystyle\int_{\mathbb{R}}K(u)\,\mathrm{d}x=\rho,\end{cases}$
where $1\leq m\in\mathbb{N}$, $\rho\in(0,+\infty)$ is a prescribed quantity, $\lambda\in\mathbb{R}$ is part of the unknown, and $F$, $G$, and $K$ are suitable functions.
To explain our motivations, let us start with taking $2G(s)=K(s)=s^{2}$. In this case, (1.1) takes the form
(1.2) $\begin{cases}\displaystyle\Bigl{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}\Bigr{)}^{m}u+\lambda u=F^{\prime}(u)\\ \displaystyle\int_{\mathbb{R}}u^{2}\,\mathrm{d}x=\rho.\end{cases}$
If $F(u)=F(|u|)$, then (1.2) arise when seeking standing-wave solutions to the Schrödinger-type evolution equation
$\mathrm{i}\frac{\partial}{\partial t}\Psi=\Bigl{(}-\frac{\partial^{2}}{\partial x^{2}}\Bigr{)}^{m}\Psi-F^{\prime}(u),$
i.e., solutions of the form $\Psi(t,x)=e^{\mathrm{i}\lambda t}u(x)$ with $u\colon\mathbb{R}\to\mathbb{R}$. Then, the $L^{2}$ constraint is justified because $|\Psi(t,\cdot)|=|u|$ for all $t\in\mathbb{R}$. Solutions to (1.2) are often referred to as normalised solutions.
A classical approach to solve (1.2) consists of finding critical points of the functional $J\colon H^{m}(\mathbb{R})\to\mathbb{R}$,
(1.3) $J(u)=\int_{\mathbb{R}}\frac{1}{2}|u^{(m)}|^{2}-F(u)\,\mathrm{d}x,$
restricted to the set
(1.4) $\mathcal{S}:=\Set{v\in H^{m}(\mathbb{R})}{\int_{\mathbb{R}}v^{2}\,\mathrm{d}x=\rho}$
under suitable conditions on $F$ that include $F^{\prime}(s)=\mathcal{O}(|s|)$ as $s\to 0$. With this approach, $-\lambda$ is nothing but the Lagrange multiplier arising from the constraint $\mathcal{S}$. Since minimisers are among the simplest examples of critical points, it makes sense to wonder whether $J|_{\mathcal{S}}$ is bounded below: this is determined by the behaviour at infinity of $F(s)$ with respect to $|s|^{2+4m}$ and, sometimes, $\rho$. In particular,
$\lim_{|s|\to+\infty}F(s)s^{-(2+4m)}\begin{cases}\leq 0\\ \in(0,+\infty)\\ =+\infty\end{cases}\Longrightarrow\quad\inf_{\mathcal{S}}J>-\infty\begin{cases}\text{for all values of }\rho\\ \text{for small values of }\rho\\ \text{for no values of }\rho\end{cases}$
(for the sake of the explanation, we assume that $\lim_{|s|\to+\infty}F(s)s^{-(2+4m)}$ exists). These three regimes are known in the literature as mass- (or $L^{2}$-) subcritical, critical, and supercritical.
It is evident, then, that the number
(1.5) $2+4m$
($2+4m/N$ in dimension $N\geq 1$) plays an important role in the geometry of $J|_{\mathcal{S}}$, which is why it is called the mass- (or $L^{2}$-) critical exponent.
When $m=1$, the mass-subcritical case, as well as the mass-critical one with $\rho$ small, were first studied by C.A. Stuart [28] and P.-L. Lions [15]; more recently, they have been dealt with, e.g., in [13, 24, 25], see also the references therein. In the mass-supercritical regime, instead, the seminal work was carried out by L. Jeanjean [10]; lately, the problem was revisited, e.g., in [11], see also the references therein. As for the mixed case, it was considered only a few years ago by N. Soave [27] and L. Jeanjean & S.-S. Lu [12].
When $m\geq 2$, instead, there is very little work: to our best knowledge, [18, 23, 29] are the only papers debating the one-dimensional case, and uniquely for $m=2$.
Going back to the article [15], it is interesting that, for $m=1$, problems more general than (1.2) are considered therein; for example, the author considers (1.1) with $pG(s)=K(s)=|s|^{p}$, $p>1$. At the same time, the recent article [14] introduced a new, non-variational method to look for solutions to (1.2) with $m=1$, which allows the mass-subcritical, -critical, and -supercritical regimes to be dealt with in the same way and where the starting point is the existence of a positive solution to
$-u^{\prime\prime}+\lambda u=F^{\prime}(u)\quad\text{in }\mathbb{R}$
with $\lambda>0$ fixed and suitable assumptions about $F$. These considerations motivates us to exploit the techniques of [14] and find solutions to (1.1) with $G$ and $K$ even more general than in [15], at least when $m=1$. When $m\geq 2$, instead, this new approach does not seem to work because of the lack of a theory about solutions to the differential equation in (1.1) with $\lambda$ fixed, and we have to rely on variational methods.
When using an approach inspired from [14], we consider the following assumptions.
  1. (A0)
    $F,G\in\mathcal{C}^{1}([0,\infty))$, $F(0)=G(0)=F^{\prime}(0)=G^{\prime}(0)=0$, $G^{\prime}(s)>0$ for all $s>0$, $\lim_{s\to 0^{+}}\frac{F(s)}{G(s)}=0$, and $\lim_{s\to+\infty}\frac{F(s)}{G(s)}=+\infty$.
  2. (A1)
    For all $s>0$ such that $F(s)>0$ there holds $Z(s):=\big{(}\frac{F}{G}\big{)}^{\prime}(s)>0$.
  3. (A2)
    $K\in\mathcal{C}([0,\infty))$, $K(0)=0$, $K(s)>0$ for all $s>0$, $s\mapsto\frac{K(s)}{\sqrt{G(s)}}$ is integrable in a right-hand neighbourhood of 00, and $\int_{0}^{\infty}\frac{K(s)}{\sqrt{G(s)}}\,\mathrm{d}s=+\infty$.
Moreover, we define
$\Phi(t):=\left(\int_{0}^{t}\frac{K(s)}{\sqrt{G(s)}}\,\mathrm{d}s\right)^{2}\ \text{for }t>0,\quad m_{0}:=\max\Set{t\geq 0}{F\leq 0\text{ on }[0,t]},$
$I_{F}:=\sqrt{2}\int_{0}^{m_{0}}\frac{K(s)}{\sqrt{|F(s)|}}\,\mathrm{d}s\in(0,+\infty],\qquad\qquad\qquad$
and, recalling the definition of $Z$ from (A1),
$L_{0}:=\limsup_{s\to 0^{+}}\frac{Z(s)}{\Phi^{\prime}(s)}\in[0,+\infty],$ $\ell_{0}:=\liminf_{s\to 0^{+}}\frac{Z(s)}{\Phi^{\prime}(s)}\in[0,+\infty],$
$L_{\infty}:=\limsup_{s\to\infty}\frac{Z(s)}{\Phi^{\prime}(s)}\in[0,+\infty],$ $\ell_{\infty}:=\liminf_{s\to\infty}\frac{Z(s)}{\Phi^{\prime}(s)}\in[0,+\infty].$
In this context, our main result reads as follows.
Theorem 1.1.
Let $m=1$ and assume that (A0)–(A2) hold. Then, a solution $(\lambda,u)\in(0,+\infty)\times\mathcal{C}^{2}(\mathbb{R})$ to (1.1) exists in each of the following cases:
  • $m_{0}=0$, $L_{0}<\ell_{\infty}$, and $\rho\in\left(\frac{\pi}{\sqrt{2\ell_{\infty}}},\frac{\pi}{\sqrt{2L_{0}}}\right)$;
  • $m_{0}=0$, $L_{\infty}<\ell_{0}$, and $\rho\in\left(\frac{\pi}{\sqrt{2\ell_{0}}},\frac{\pi}{\sqrt{2L_{\infty}}}\right)$;
  • $m_{0}>0$, $I_{F}>\frac{\pi}{\sqrt{2\ell_{\infty}}}$, and $\rho\in\left(\frac{\pi}{\sqrt{2\ell_{\infty}}},I_{F}\right)$;
  • $m_{0}>0$, $F^{\prime}(m_{0})\neq 0$, $I_{F}<\frac{\pi}{\sqrt{2L_{\infty}}}$, and $\rho\in\left(I_{F},\frac{\pi}{\sqrt{2L_{\infty}}}\right)$.
Furthermore, $u^{\prime}\in L^{2}(\mathbb{R})$.
As examples for such $G$ and $K$ we propose
$G(s)=\frac{1}{p}s^{p}\quad\text{and}\quad K(s)=s^{q},$
with $p>1$ and $q>\max\{p/2-1,0\}$ (observe that this includes the case $p=q=2$). Then, from Theorem 1.1 we obtain immediately the following outcome.
Corollary 1.2.
Let $m=1$, $p>1$, $q>\max\{p/2-1,0\}$, and assume that
  1. (a0)
    $F\in\mathcal{C}^{1}([0,\infty))$, $\lim_{s\to 0^{+}}\frac{F(s)}{s^{p}}=0$, and $\lim_{s\to+\infty}\frac{F(s)}{s^{p}}=+\infty$.
  2. (a1)
    For all $s>0$ such that $F(s)>0$ there holds $F^{\prime}(s)s-pF(s)>0$.
For $s>0$, define the quantities
$K_{0}:=\limsup_{s\to 0^{+}}\frac{F^{\prime}(s)s-pF(s)}{s^{2q+2}},$ $k_{0}:=\liminf_{s\to 0^{+}}\frac{F^{\prime}(s)s-pF(s)}{s^{2q+2}},$
$K_{\infty}:=\limsup_{s\to\infty}\frac{F^{\prime}(s)s-pF(s)}{s^{2q+2}},$ $k_{\infty}:=\liminf_{s\to\infty}\frac{F^{\prime}(s)s-pF(s)}{s^{2q+2}}.$
Then, a solution $(\lambda,u)\in(0,+\infty)\times\mathcal{C}^{2}(\mathbb{R})$ to (1.1) exists in every of the following cases:
  • $m_{0}=0$, $K_{0}<k_{\infty}$, and $\rho\in\left(\frac{\pi}{\sqrt{(q-p/2+1)k_{\infty}}},\frac{\pi}{\sqrt{(q-p/2+1)K_{0}}}\right)$;
  • $m_{0}=0$, $K_{\infty}<k_{0}$, and $\rho\in\left(\frac{\pi}{\sqrt{(q-p/2+1)k_{0}}},\frac{\pi}{\sqrt{(q-p/2+1)K_{\infty}}}\right)$;
  • $m_{0}>0$, $I_{F}<\frac{\pi}{\sqrt{(q-p/2+1)K_{\infty}}}$, and $\rho\in\left(\frac{\pi}{\sqrt{(q-p/2+1)k_{\infty}}},I_{F}\right)$;
  • $m_{0}>0$, $F^{\prime}(m_{0})\neq 0$, $I_{F}<\frac{\pi}{\sqrt{(q-p/2+1)K_{\infty}}}$, and $\rho\in\left(I_{F},\frac{\pi}{\sqrt{(q-p/2+1)K_{\infty}}}\right)$.
Furthermore, $u^{\prime}\in L^{2}(\mathbb{R})$.
Remark 1.3.
Here are some observations about the case $p=q=2$ in Corollary 1.2.
  • (i)
    The number $62q+2=$ 6, which appears in the definition of $K_{0}$, $k_{0}$, $K_{\infty}$, and $k_{\infty}$, is exactly the exponent (1.5) with $m=1$.
  • (ii)
    At first glance, when $F>0$ on $(0,+\infty)$, Corollary 1.2 is weaker than [14, Theorem 1.2] in dimension $1$ (see also [14, Remark 2.6]) because of (a1); however, such an assumption is needed (cf. Remark 2.1 below), hence [14, Theorem 6.6], which [14, Theorem 1.2] is based on, contains a (small) gap.
  • (iii)
    Since we can admit sign-changing terms $F$, Corollary 1.2 improves the one-dimensional case of [14, Theorem 1.2]. Additionally, it extends the one-dimensional existence results in [27] to the case of non-linearities more general than the sum of two powers.
Now, we turn to the case where $m$ is any positive integer. We begin by considering assumptions that describe the mass-subcritical and -critical cases.
  1. (f0)
    $F\in\mathcal{C}^{1}(\mathbb{R})$ and $F^{\prime}(s)=\mathcal{O}(|s|)$ as $s\to 0$.
  2. (f1)
    $\lim_{s\to 0}F(s)s^{-2}=0$.
  3. (f2)
    $\sigma:=\limsup_{|s|\to+\infty}F(s)s^{-(2+4m)}<+\infty$.
  4. (f3)
    $\lim_{s\to 0}F(s)s^{-(2+4m)}=+\infty$.
We recall from (1.3) and (1.4) the definitions of $J$ and $\mathcal{S}$ and introduce the set
$\mathcal{D}:=\Set{v\in H^{m}(\mathbb{R})}{\int_{\mathbb{R}}v^{2}\,\mathrm{d}x\leq\rho},$
which was first used in [4] and then exploited, e.g., in [3, 5, 6, 17, 20, 21, 24, 16]. Moreover, let us recall the Gagliardo–Niremberg inequality [7, 8, 22], here expressed in the one-dimensional case: for every $p>2$ there exists $C_{p}>0$ such that for all $v\in H^{m}(\mathbb{R})$ there holds
(1.6) $|v|_{p}\leq C_{p}|v^{(m)}|_{2}^{\delta_{p}}|v|_{2}^{1-\delta_{p}},$
and $C_{p}$ is sharp, where $\delta_{p}=(1/2-1/p)/m$ and $|\cdot|_{q}$ denotes the norm in $L^{q}(\mathbb{R})$, $q\in[1,+\infty]$.
Our existence result in this regime is the following.
Theorem 1.4.
If (f0)–(f3) are satisfied and $12\sigma C_{2+4m}^{2+4m}\rho^{2m}<$ 1, then there exist $u\in\mathcal{S}$ and $\lambda>0$ such that $J(u)=\min_{\mathcal{D}}J<0$ and $(\lambda,u)$ is a solution to (1.2).
The proof of Theorem 1.4 follows verbatim that of [24, Theorem 1.1] once proved that every solution $u\in H^{m}(\mathbb{R})$ to the differential equation in (1.2) with $\lambda\in\mathbb{R}$ fixed satisfies the Pohožaev identity if (f0) holds, which is done in Proposition 3.1 below. For this reason, we omit it.
Next, we move to the mass-supercritical case. Let us define $H(s):=F^{\prime}(s)s-2F(s)$ for $s\in\mathbb{R}$. We assume the following conditions.
  1. (F0)
    $F,H\in\mathcal{C}^{1}(\mathbb{R})$ and $|F^{\prime}(s)|+|H^{\prime}(s)|=\mathcal{O}(|s|)$ as $s\to 0$.
  2. (F1)
    $\eta:=\limsup_{s\to 0}H(s)s^{-(2+4m)}<+\infty$.
  3. (F2)
    $\lim_{|s|\to+\infty}F(s)s^{-(2+4m)}=+\infty$
  4. (F3)
    $(2+4m)H(s)\leq H^{\prime}(s)s$ for all $s\in\mathbb{R}$.
  5. (F4)
    $0\leq 4mF(s)\leq H(s)$ for all $s\in\mathbb{R}$.
Since $J|_{\mathcal{S}}$ is unbounded below under (F0)–(F4), cf. the proof of Lemma 3.2 below, we follow the approach in [4], where – moreover – examples of such functions $F$ can be found.
Let us recall the Nehari and Pohožaev identities associated with (1.2), i.e.,
$\int_{\mathbb{R}}|u^{(m)}|^{2}+\lambda u^{2}\,\mathrm{d}x$ $=\int_{\mathbb{R}}F^{\prime}(u)u\,\mathrm{d}x$
$\int_{\mathbb{R}}(1-2m)|u^{(m)}|^{2}+\lambda u^{2}\,\mathrm{d}x$ $=2\int_{\mathbb{R}}F(u)\,\mathrm{d}x$
respectively. Then, every $u\in H^{m}(\mathbb{R})\setminus\{0\}$ that solves the differential equation in (1.2) for some $\lambda\in\mathbb{R}$ belongs to the set
:=$\mathcal{M}:=\Set{v\in H^{m}(\mathbb{R})\setminus\{0\}}{\int_{\mathbb{R}}|v^{(m)}|^{2}\,\mathrm{d}x=\frac{1}{2m}\int_{\mathbb{R}}H(v)\,\mathrm{d}x}.$
If $H\in\mathcal{C}^{1}(\mathbb{R})$, $H^{\prime}(s)=\mathcal{O}(|s|)$ as $s\to 0$, and $H(\xi_{0})>0$ for some $\xi_{0}\neq 0$, one easily proves that $\mathcal{M}$ is a manifold of class $\mathcal{C}^{1}$ and co-dimension $1$; see, e.g., [5, Lemma 4.1].
We consider the following condition, which will be paired with (F3):
(1.7) $\int_{\mathbb{R}}H^{\prime}(u)u-(2+4m)H(u)\,\mathrm{d}x>0\quad\forall u\in H^{m}(\mathbb{R})\setminus\{0\}.$
Note that (1.7) is satisfied if and only if $H(s)s-(2+4m)H(s)\geq 0$ holds for every $s\in\mathbb{R}$ and the strict inequality holds along two sequences $s^{\prime}_{n}\to 0^{+}$ and $s^{\prime\prime}_{n}\to 0^{-}$ (cf. [4, Lemma 2.1]).
Finally, we introduce the condition
(1.8) $\eta C_{2+4m}^{2+4m}\rho^{2m}<2m,$
which allows us to deal with a non-linearity with mass-critical growth at the origin.
Our existence result in this regime is the following.
Theorem 1.5.
If (F0)–(F4) and (1.8) hold, then there exists $u\in\mathcal{D}\cap\mathcal{M}$ such that $J(u)=\min_{\mathcal{D}\cap\mathcal{M}}J>0$. If, moreover, (1.7) holds, then there exists $\lambda>0$ such that $(\lambda,u)$ is solution to (1.1) – in particular, $u\in\mathcal{S}$.
The proof of Theorem 1.5 is modelled on that of [5, Theorem 3.3]; nonetheless, since the one-dimensional setting requires some modifications, we provide it in Section 3 for the reader’s convenience. Theorem 1.1, instead, is proved in Section 2.
2. The global-branch approach
Throughout this section, $m=1$, and we assume (A0)–(A2).
In view of (A0), for every $\lambda>0$ there exists $m_{\lambda}\in(0,+\infty)$ such that $W_{\lambda}:=\lambda G-F$ is positive on $(0,m_{\lambda})$ and $W_{\lambda}(0)=W^{\prime}_{\lambda}(0)=W_{\lambda}(m_{\lambda})=0$. On the other hand, (A1) implies that $W^{\prime}_{\lambda}(m_{\lambda})=-Z(m_{\lambda})G(m_{\lambda})<0.$ Consequently, for every $\lambda>0$, there exists a homoclinic orbit $u_{\lambda}$ corresponding to $W_{\lambda}$ (cf. for instance [2, Theorem 5] or [1, Theorem 5.4] [^1] Both these theorems require additional regularity for $F$ and $G$ (at least $\mathcal{C}_{\textup{loc}}^{1,1}$); however, similar arguments as in [1] can be repeated with minor modifications when $F$ and $G$ are merely $\mathcal{C}^{1}$.) satisfying the following properties:
  • $u_{\lambda}\in\mathcal{C}^{2}(\mathbb{R})$ is a non-negative and even solution of $-u^{\prime\prime}+\lambda G^{\prime}(u)=F^{\prime}(u)$.
  • $\lim_{|x|\to\infty}u_{\lambda}(x)=0$ and $m_{\lambda}=u_{\lambda}(0)=\max_{\mathbb{R}}u_{\lambda}$.
  • $|u^{\prime}_{\lambda}(x)|^{2}=2W_{\lambda}(u_{\lambda}(x))$ for all $x\in\mathbb{R}$ (equipartition relation).
  • Setting $T_{\lambda}:=\int_{0}^{m_{\lambda}}\frac{\mathrm{d}u}{\sqrt{2W_{\lambda}(u)}}\in(0,+\infty]$ [^2] Observe that $T_{\lambda}=+\infty$ if $W_{\lambda}\in\mathcal{C}^{1,1}([0,\varepsilon])$ for some $\varepsilon>0$., $u_{\lambda}$ is increasing on $(-T_{\lambda},0)$ and decreasing on $(0,T_{\lambda})$, while $u_{\lambda}(x)=0$, if $|x|\geq T_{\lambda}$.
  • $u^{\prime}_{\lambda}\in L^{2}(\mathbb{R})$.
Remark 2.1.
We point out that the assumption $F^{\prime}(t)>0$ for all $t>0$ is not sufficient to ensure, for every $\lambda>0$, the existence of the homoclinic orbit $u_{\lambda}$, since we may have $W^{\prime}_{\lambda}(m_{\lambda})=0$ for some $\lambda>0$. For example, when $G(s)=s^{2}/2$, taking
$F(s)=\begin{cases}\frac{1}{2}s^{2}+\cos(s)-1&\text{if }s\in[0,2\pi]\\ \frac{1}{2}s^{2}+(s-2\pi)^{p}&\text{if }s>2\pi\end{cases}$
with $p>2$ we see that $m_{1}=2\pi$ and $W_{1}^{\prime}(m_{1})=0$, thus no non-trivial non-negative solutions to $-u^{\prime\prime}+u=F^{\prime}(u)$ that vanish at infinity exist in view of [2, Theorem 5].
Recalling that $m_{0}=\max\{t\geq 0:F\leq 0\text{ on }[0,t]\}$, we have $F(m_{0})=0$, and in view of (A1), $F$ and $F^{\prime}$ are positive on $(m_{0},+\infty)$. In addition, the following properties hold.
Lemma 2.2.
The function $(0,\infty)\ni\lambda\mapsto m_{\lambda}\in(0,\infty)$ is increasing, continuous, $\lim_{\lambda\to+\infty}m_{\lambda}=+\infty$, and $\lim_{\lambda\to 0^{+}}m_{\lambda}=m_{0}$.
Proof.
Observe that (A0) implies that $G(s)>0$ if $s>0$. Let $0<\lambda<\Lambda<\infty$ and $s\in(0,m_{\lambda}]$; then,
$W_{\Lambda}(s)=\Lambda G(s)-F(s)>\lambda G(s)-F(s)=W_{\lambda}(s)\geq 0,$
which shows that $m_{\Lambda}>m_{\lambda}$.
The continuity follows from the implicit function theorem applied to the function $(\lambda,s)\mapsto W_{\lambda}(s)$ because, as remarked above, $W^{\prime}_{\lambda}(m_{\lambda})<0$.
Next, observe that the two limits in the statement exist in virtue of the monotonicity of $\lambda\mapsto m_{\lambda}$. Assume by contradiction that $\lim_{\lambda\to+\infty}m_{\lambda}=:M<+\infty$. Then
$0=\lim_{\lambda\to+\infty}W_{\lambda}(m_{\lambda})=\lim_{\lambda\to+\infty}\lambda G(m_{\lambda})-F(m_{\lambda})=+\infty,$
which is impossible.
Finally, observe that the definition of $m_{0}$ and the fact that $F(m_{\lambda})=m_{\lambda}G(m_{\lambda})>0$ imply that $m_{0}<m_{\lambda}$. Denoting $M:=\lim_{\lambda\to 0^{+}}m_{\lambda}\geq m_{0}$, we have
$0=\lim_{\lambda\to 0^{+}}W_{\lambda}(m_{\lambda})=\lim_{\lambda\to 0^{+}}\lambda G(m_{\lambda})-F(m_{\lambda})=F(M),$
and the statement follows from the fact that $F$ is positive on $(m_{0},+\infty)$. ∎
Lemma 2.3.
For every $\lambda>0$ there holds
$\rho_{\lambda}:=\int_{\mathbb{R}}K(u_{\lambda}(x))\,\mathrm{d}x=\sqrt{2}\int_{0}^{m_{\lambda}}\frac{K(u)}{\sqrt{G(u)}}\left(\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\right)^{-1/2}\,\mathrm{d}u<\infty.$
Furthermore, $\lambda\mapsto\rho_{\lambda}$ is continuous.
Proof.
Observe that $u_{\lambda}|_{(-T_{\lambda},0)}\colon(-T_{\lambda},0)\to(0,m_{\lambda})$ is a diffeomorphism and, from the equipartition relation, $u_{\lambda}^{\prime}(x)=\sqrt{2W_{\lambda}(u_{\lambda}(x))}$. Consequently,
(2.1) $\rho_{\lambda}=\int_{\mathbb{R}}K(u_{\lambda}(x))\,\mathrm{d}x=2\int_{-T_{\lambda}}^{0}K(u_{\lambda}(x))\,\mathrm{d}x=\sqrt{2}\int_{0}^{m_{\lambda}}\frac{K(u)}{\sqrt{W_{\lambda}(u)}}\,\mathrm{d}u.$
Since, from (A0) and (A2), $F(s)/G(s)\to 0$ as $s\to 0^{+}$ and $K/\sqrt{G}$ is integrable in a right-hand neighbourhood of 00, we have that $K/\sqrt{W_{\lambda}}$ is integrable in a right-hand neighbourhood of 00 as well. This, together with the property that $W_{\lambda}^{\prime}(m_{\lambda})\neq 0$, yields that $\rho_{\lambda}<+\infty$. Moreover, the continuity of $\lambda\mapsto\rho_{\lambda}$ follows from the one of $\lambda\mapsto m_{\lambda}$ (Lemma 2.2) and the facts above. Finally,
(2.2) $\begin{split}\int_{0}^{m_{\lambda}}\frac{K(u)}{\sqrt{W_{\lambda}(u)}}\,\mathrm{d}u&=\int_{0}^{m_{\lambda}}\frac{K(u)}{\sqrt{G(u)}}\left(\lambda-\frac{F(u)}{G(u)}\right)^{-1/2}\,\mathrm{d}u\\ &=\int_{0}^{m_{\lambda}}\frac{K(u)}{\sqrt{G(u)}}\left(\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\right)^{-1/2}\,\mathrm{d}u,\end{split}$
and the conclusion follows from (2.1)–(2.2). ∎
Next, we determine the asymptotic behaviour of $\rho_{\lambda}$ as $\lambda\to 0^{+}$ and $\lambda\to\infty$.
Lemma 2.4.
Assume (A0)–(A2). With the convention that $\frac{1}{0}=+\infty$ and $\frac{1}{+\infty}=0$, the following holds.
  1. (1)
    If $m_{0}=0$, then $\liminf_{\lambda\to 0^{+}}\rho_{\lambda}\geq\frac{\pi}{\sqrt{2L_{0}}}$ and $\limsup_{\lambda\to 0^{+}}\rho_{\lambda}\leq\frac{\pi}{\sqrt{2\ell_{0}}}$.
  2. (2)
    If $m_{0}>0$, then $\liminf_{\lambda\to 0^{+}}\rho_{\lambda}\geq I_{F}$. If, in addition, $F^{\prime}(m_{0})\neq 0$, then $\lim_{\lambda\to 0^{+}}\rho_{\lambda}=I_{F}$.
  3. (3)
    $\liminf_{\lambda\to+\infty}\rho_{\lambda}\geq\frac{\pi}{\sqrt{2L_{\infty}}}$ and $\limsup_{\lambda\to+\infty}\rho_{\lambda}\leq\frac{\pi}{\sqrt{2\ell_{\infty}}}$.
Proof.
(1) Given $L\in(L_{0},+\infty)$, there exists $\delta_{L}>0$ such that $Z(s)\leq L\Phi^{\prime}(s)$ for all $s\in(0,\delta_{L})$. In addition, from Lemma 2.2, $m_{\lambda}<\delta_{L}$ holds for $\lambda\in(0,\Lambda_{L})$, with $\Lambda_{L}>0$ small enough. Therefore, we have for $\lambda\in(0,\Lambda_{L})$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\leq L\bigl{(}\Phi(m_{\lambda})-\Phi(u)\bigr{)}\quad\forall u\in[0,m_{\lambda}]$
and, using Lemma 2.3 and the change of variable $t=\sqrt{\frac{\Phi(u)}{\Phi(m_{\lambda})}}$,
$\rho_{\lambda}\geq\sqrt{\frac{2}{L}}\int_{0}^{m_{\lambda}}\frac{K(u)(G(u))^{-1/2}}{(\Phi(m_{\lambda})-\Phi(u))^{1/2}}\,\mathrm{d}u=\sqrt{\frac{2}{L}}\int_{0}^{1}\frac{1}{(1-t^{2})^{1/2}}\,\mathrm{d}t=\frac{\pi}{\sqrt{2L}},$
which proves the first part of (1). The second one is proved similarly.
(2) From (2.1) and the monotone convergence theorem,
$\liminf_{\lambda\to 0^{+}}\rho_{\lambda}\geq\lim_{\lambda\to 0^{+}}\sqrt{2}\int_{0}^{m_{0}}\frac{K(u)}{\sqrt{W_{\lambda}(u)}}\,\mathrm{d}u=\sqrt{2}\int_{0}^{m_{0}}\frac{K(u)}{\sqrt{|F(u)|}}\,\mathrm{d}u=I_{F}.$
Next, if $F^{\prime}(m_{0})\neq 0$, which implies $Z(m_{0})\neq 0$, then – cf. (2.2)
$\int_{m_{0}}^{m_{\lambda}}\frac{K(u)}{\sqrt{W_{\lambda}(u)}}\,\mathrm{d}u$ $=\int_{m_{0}}^{m_{\lambda}}\frac{K(u)(G(u))^{-1/2}}{\sqrt{\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}}}\,\mathrm{d}u=\int_{m_{0}}^{m_{\lambda}}\frac{K(u)(G(u))^{-1/2}}{\sqrt{\int_{u}^{m_{\lambda}}Z(s)\,\mathrm{d}s}}\,\mathrm{d}u$
$=\mathcal{O}(\sqrt{m_{\lambda}-m_{0}})\to 0\quad\text{as }\lambda\to 0^{+},$
therefore $\lim_{\lambda\to 0^{+}}\rho_{\lambda}=I_{F}$.
(3) Let us begin with the first part. Given $L\in(L_{\infty},+\infty)$, there exists $S_{L}>m_{0}$ such that $Z(s)\leq L\Phi^{\prime}(s)$ for all $s\geq S_{L}$. In addition, from Lemma 2.2, $m_{\lambda}>S_{L}$ holds for $\lambda\in(\Lambda_{L},\infty)$, with $\Lambda_{L}>0$ large enough. Therefore, we have for $\lambda\in(\Lambda_{L},\infty)$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\leq L\Phi(m_{\lambda})-\Phi(u)\quad\text{for all }u\in[S_{L},m_{\lambda}].$
This, together with (A2) and Lemmas 2.2 and 2.3, implies
$\rho_{\lambda}$ $\geq\sqrt{2}\int_{S_{L}}^{m_{\lambda}}\frac{K(u)}{\sqrt{G(u)}}\left(\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}\right)^{-1/2}\,\mathrm{d}u$
$\geq\sqrt{\frac{2}{L}}\int_{S_{L}}^{m_{\lambda}}\frac{K(u)(G(u))^{-1/2}}{(\Phi(m_{\lambda})-\Phi(u))^{1/2}}\,\mathrm{d}u=\sqrt{\frac{2}{L}}\int_{\sqrt{\Phi(S_{L})/\Phi(m_{\lambda})}}^{1}\frac{1}{(1-t^{2})^{1/2}}\,\mathrm{d}t$
$\to\sqrt{\frac{2}{L}}\int_{0}^{1}\frac{1}{(1-t^{2})^{1/2}}\,\mathrm{d}t=\frac{\pi}{\sqrt{2L}}\quad\text{as }\lambda\to+\infty.$
Now we move to the second part. Given $L\in(0,\ell_{\infty})$, there exists $S_{L}>m_{0}$ such that $Z(s)\geq L\Phi^{\prime}(s)$ for all $s\geq S_{L}$. Additionally, since $F/G$ is non-positive on $[0,m_{0}]$ and, from (A1), positive and increasing on $(m_{0},+\infty)$, there holds
$\frac{F(S_{L})}{G(S_{L})}\geq\frac{F(u)}{G(u)}\quad\text{for all }u\in(0,S_{L}].$
In addition, from Lemma 2.2, $m_{\lambda}>S_{L}$ holds for $\lambda\in(\Lambda_{L},\infty)$, with $\Lambda_{L}>0$ large enough. Therefore, we have for $\lambda\in(\Lambda_{L},\infty)$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}$ $\geq L\Phi(m_{\lambda})-\Phi(u)\quad\text{for all }u\in[S_{L},m_{\lambda}],$
$\frac{F(m_{\lambda})}{G(m_{\lambda})}-\frac{F(u)}{G(u)}$ $\geq L\Phi(m_{\lambda})-\Phi(S_{L})\quad\text{for all }u\in(0,S_{L}],$
This, together with Lemma 2.3, implies
$\rho_{\lambda}\leq\sqrt{\frac{2}{L}}\bigl{(}I_{1}(\lambda)+I_{2}(\lambda)\bigl{)}\quad\text{for all }\lambda\in(\Lambda_{L},\infty),$
where, thanks to (A2),
$I_{1}(\lambda)$ $=\int_{0}^{S_{L}}\frac{K(u)(G(u))^{-1/2}}{(\Phi(m_{\lambda})-\Phi(S_{L}))^{1/2}}\,\mathrm{d}u\to 0\quad\text{as }\lambda\to+\infty,$
$I_{2}(\lambda)$ $=\int_{S_{L}}^{m_{\lambda}}\frac{K(u)(G(u))^{-1/2}}{(\Phi(m_{\lambda})-\Phi(u))^{1/2}}\,\mathrm{d}u\leq\int_{0}^{m_{\lambda}}\frac{K(u)(G(u))^{-1/2}}{(\Phi(m_{\lambda})-\Phi(u))^{1/2}}\,\mathrm{d}u=\frac{\pi}{2},$
which proves the second part of (3). ∎
Proof of Theorem 1.1.
It follows from Lemmas 2.3 and 2.4. ∎
Remark 2.5.
Similarly, non-existence results for problem (1.1) can be established in view of Lemma 2.4 and the continuity of the function $\lambda\mapsto\rho_{\lambda}$.
3. The poly-harmonic case
We begin by proving that every solution to the differential equation in (1.2) satisfies the Pohožaev identity. In fact, we can state a more general result.
Proposition 3.1 (Pohožaev identity).
Let $N\in[1,2m-1]$ be an integer, $g\in\mathcal{C}(\mathbb{R})$ such that $g(s)=\mathcal{O}(|u|)$ as $s\to 0$, and define $G(s):=\int_{0}^{s}g(t)\,\mathrm{d}t$. If $u\in H^{m}(\mathbb{R})$ is a weak solution to
(3.1) $(-\Delta)^{m}u=g(u)\quad\text{in }\mathbb{R}^{N},$
then
$(N-2m)\int_{\mathbb{R}^{N}}|\nabla^{m}u|^{2}\,\mathrm{d}x=2N\int_{\mathbb{R}^{N}}G(u)\,\mathrm{d}x.,$
where
$\nabla^{m}u$ :=$:=\begin{cases}\Delta^{m/2}u&\text{if }m\text{ is even,}\\ \nabla\Delta^{(m-1)/2}u&\text{if }m\text{ is odd.}\end{cases}$
Proof.
Since, from the Sobolev embedding, $|u|_{\infty}<+\infty$, we have that $|g(u)|\lesssim|u|$ a.e. in $\mathbb{R}^{N}$, hence $g\circ u\in L^{2}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$. Then, from [26, Lemma 3.1], $u\in W^{2m,p}_{\textup{loc}}(\mathbb{R}^{N})$ for every $p\in[1,+\infty)$. Now, one can follow [5, Proof of Proposition 2.5]. We sketch the proof for the reader’s convenience.
For every $n\geq 1$, let $\psi_{n}\in\mathcal{C}^{1}_{0}(\mathbb{R}^{N})$ radially symmetric such that $0\leq\psi_{n}\leq 1$, $\psi_{n}(x)=1$ for every $|x|\leq n$, $\psi_{n}(x)=0$ for every $|x|\geq 2n$, and $|x||\nabla\psi_{n}(x)|\lesssim 1$ for every $x\in\mathbb{R}^{N}$.
Next, observe that the following identities hold true:
$g(u)(\nabla u\cdot x)\psi_{n}=$ $\;\nabla\cdot\bigl{(}\psi_{n}G(u)x\bigr{)}-N\psi_{n}G(u)-G(u)\nabla\psi_{n}\cdot x,$
$\Delta^{2k+1}u(\nabla u\cdot x)\psi_{n}=$ $\;\nabla\cdot\biggl{[}\biggl{(}\Delta^{k}(x\cdot\nabla u)\nabla\Delta^{k}u-\frac{|\nabla\Delta^{k}u|^{2}}{2}x$
$-\sum_{j=0}^{k-1}\Delta^{2k-j}u\nabla\Delta^{j}(\nabla u\cdot x)+\sum_{j=0}^{k-1}\Delta^{j}(\nabla u\cdot x)\nabla\Delta^{2k-j}u\biggr{)}\psi_{n}\biggr{]}$
$+\frac{N-4k-2}{2}|\nabla\Delta^{k}u|^{2}\psi_{n}-\biggl{(}\Delta^{k}(\nabla u\cdot x)\nabla\Delta^{k}u-\frac{|\nabla\Delta^{k}u|^{2}}{2}x$
$-\sum_{j=0}^{k-1}\Delta^{2k-j}u\nabla\Delta^{j}(\nabla u\cdot x)+\sum_{j=0}^{k-1}\Delta^{j}(\nabla u\cdot x)\nabla\Delta^{2k-j}u\biggr{)}\cdot\nabla\psi_{n},$
$\Delta^{2k}u(\nabla u\cdot x)\psi_{n}=$ $\;\nabla\cdot\biggl{[}\biggl{(}\frac{1}{2}(\Delta^{k}u)^{2}x+(\nabla u\cdot x)\nabla\Delta^{2k-1}u$
$+\sum_{j=0}^{k-2}\Delta^{j+1}(\nabla u\cdot x)\nabla\Delta^{2k-j-2}u$
$-\sum_{j=0}^{k-1}\Delta^{2k-j-1}u\nabla\Delta^{j}(\nabla u\cdot x)\biggr{)}\psi_{n}\biggr{]}+\frac{4k-N}{2}(\Delta^{k}u)^{2}\psi_{n}$
$-\biggl{(}\frac{1}{2}(\Delta^{k}u)^{2}x+(\nabla u\cdot x)\nabla\Delta^{2k-1}u$
$+\sum_{j=0}^{k-2}\Delta^{j+1}(\nabla u\cdot x)\nabla\Delta^{2k-j-2}u$
$-\sum_{j=0}^{k-1}\Delta^{2k-j-1}u\nabla\Delta^{j}(\nabla u\cdot x)\biggr{)}\cdot\nabla\psi_{n}.$
Multiplying both sides of (3.1) by $\psi_{n}\nabla u\cdot x$, using the identities above, and integrating over $\mathbb{R}^{N}$, we obtain
(3.2) $\begin{split}0&=\int_{\mathbb{R}^{N}}\left(-(-\Delta)^{m}u+g(u)\right)\psi_{n}\nabla u\cdot x\,\mathrm{d}x\\ &=\int_{\mathbb{R}^{N}}\frac{1}{2}|\nabla^{m}u|^{2}\nabla\psi_{n}\cdot x+\mathcal{X}\cdot\nabla\psi_{n}+\frac{N-2m}{2}\psi_{n}|\nabla^{m}u|^{2}-N\psi_{n}G(u)\\ &\qquad-G(u)\nabla\psi_{n}\cdot x+\nabla\cdot\left[\psi_{n}\left(-\mathcal{X}-\frac{1}{2}|\nabla^{m}u|^{2}x+G(u)x\right)\right]\,\mathrm{d}x,\end{split}$
where
$\mathcal{X}:=\begin{cases}\displaystyle-\Delta^{k}(\nabla u\cdot x)\nabla\Delta^{k}u+\sum_{j=0}^{k-1}\Delta^{2k-j}u\nabla\Delta^{j}(\nabla u\cdot x)-\sum_{j=0}^{k-1}\Delta^{j}(\nabla u\cdot x)\nabla\Delta^{2k-j}u&\\ \displaystyle\nabla u\cdot x\nabla\Delta^{2k-1}u+\sum_{j=0}^{k-2}\Delta^{j+1}(\nabla u\cdot x)\nabla\Delta^{2k-j-2}u-\sum_{j=0}^{k-1}\Delta^{2k-j-1}u\nabla\Delta^{j}(\nabla u\cdot x)\end{cases}$
if $m=2k+1$ or $m=2k$ respectively.
Finally, from the properties of $\psi_{n}$ and the dominated convergence theorem, we conclude the proof letting $n\to+\infty$ in (3.2). ∎
For $u\in H^{m}(\mathbb{R})\setminus\{0\}$ and $s>0$, let us define $s\star u:=\sqrt{s}\,u(s\cdot)$ and $\varphi_{u}(s):=J(s\star u)$. Note that $|s\star u|_{2}=|u|_{2}$ and that $s\star u\in\mathcal{M}$ if and only if $\varphi_{u}^{\prime}(s)=0$.
Lemma 3.2.
Assume (F0)–(F4) and (1.8) hold. For every $u\in H^{m}(\mathbb{R})$ such that
(3.3) $\eta|u|_{2+4m}^{2+4m}<2m|u^{(m)}|_{2}^{2}$
there exist $0<a\leq b<+\infty$ such that $\varphi_{u}$ is increasing on $(0,a)$, decreasing on $(b,+\infty)$, and $\varphi_{u}\equiv\max\varphi_{u}$ on $[a,b]$. If, moreover, (1.7) is satisfied, then $a=b$.
Notice that every $u\in\mathcal{D}\cap H^{m}(\mathbb{R})\setminus\{0\}$ satisfies (3.3) if (1.8) holds.
Proof.
Let $u\in H^{m}(\mathbb{R})$ as in the assumptions. From (F1), (F4), and the continuity of $F$, there exists $c>0$ such that
$|F(t)|\leq ct^{2+4m}\quad\forall t\in\overline{B(0,|u|_{\infty})},$
whence
$\varphi_{u}(s)=\frac{s^{2m}}{2}|u^{(m)}|_{2}^{2}-\int_{\mathbb{R}}\frac{F(s^{1/2}u)}{s}\,\mathrm{d}x\to 0$
as $s\to 0^{+}$. Moreover,
$\frac{\varphi_{u}(s)}{s^{2m}}=\frac{|u^{(m)}|_{2}^{2}}{2}-\int_{\mathbb{R}}\frac{F(s^{1/2}u)}{s^{1+2m}}\,\mathrm{d}x,$
with $\lim_{s\to+\infty}\int_{\mathbb{R}}F(s^{1/2}u)/s^{1+2m}\,\mathrm{d}x=+\infty$ from (F2) and Fatou’s Lemma. This proves that $\lim_{s\to+\infty}\varphi_{u}(s)=-\infty$.
Now, fix $\varepsilon>0$ such that $(\eta+2m\varepsilon)|u|_{2+4m}^{2+4m}<2m|u^{(m)}|_{2}^{2}$. From (F1), (F4), and the continuity of $F$, there exists $C=C(\varepsilon,|u|_{\infty})>0$ such that
$F(t)\leq\left(\frac{\eta}{4m}+\varepsilon\right)t^{2+4m}+Ct^{4+4m}\quad\forall t\in\overline{B(0,|u|_{\infty})},$
whence, using also (3.3),
$\varphi_{u}(s)$ $\geq\frac{s^{2m}}{2}|u^{(m)}|_{2}^{2}-\frac{1}{s}\left[\left(\frac{\eta}{4m}+\varepsilon\right)|s^{1/2}u|_{2+4m}^{2+4m}+C|s^{1/2}u|_{4+4m}^{4+4m}\right]$
$=\frac{s^{2m}}{2}\left[|u^{(m)}|_{2}^{2}-\left(\frac{\eta}{2m}+\varepsilon\right)|u|_{2+4m}^{2+4m}\right]-C|u|_{4+4m}^{4+4m}s^{2m+1},$
which proves that $\varphi_{u}(s)>0$ if $s\ll 1$. So far we have proved that $\varphi_{u}$ attains its positive maximum. To conclude, note that
$\varphi_{u}^{\prime}(s)=ms^{2m-1}\left(|u^{(m)}|_{2}^{2}-\frac{1}{2m}\int_{\mathbb{R}}\frac{H(s^{1/2}u)}{s^{1+2m}}\,\mathrm{d}x\right),$
where the function
$s\mapsto\int_{\mathbb{R}}\frac{H(s^{1/2}u)}{s^{1+2m}}\,\mathrm{d}x$
is non-decreasing from (F0) and (F3), and increasing if (1.7) holds. ∎
Remark 3.3.
If (F0)–(F4) and (1.8) hold, then given $u\in\mathcal{S}\cap H^{m}(\mathbb{R})\setminus\{0\}$, it follows from the proof of Lemma 3.2 that $\varphi_{u}^{\prime}(s)=0$ for some $s>0$. Thus, we have $s\star u\in\mathcal{S}\cap\mathcal{M}$, and $\mathcal{S}\cap\mathcal{M}\neq\emptyset$.
Remark 3.4.
If (F0), (F3), and (1.7) hold, then $\Set{u\in\mathcal{M}}{\varphi_{u}^{\prime\prime}(1)=0}=\emptyset$. As a matter of fact, from $\varphi_{u}^{\prime}(1)=\varphi_{u}^{\prime\prime}(1)=0$ we obtain
$\int_{\mathbb{R}}(2+4m)H(u)-H^{\prime}(u)u\,\mathrm{d}x=0,$
which contradicts (1.7) because $u\neq 0$.
Lemma 3.5.
If (F0), (F1), and (1.8) hold, then $\inf_{u\in\mathcal{D}\cap\mathcal{M}}|u^{(m)}|_{2}>0$.
Proof.
Let $u\in\mathcal{D}\cap\mathcal{M}$. If $\|u\|_{H^{m}}>\sqrt{2\rho}$, then
$|u^{(m)}|_{2}^{2}=\|u\|_{H^{m}}^{2}-|u|_{2}^{2}>2\rho-\rho=\rho,$
hence we assume that $\|u\|_{H^{m}}\leq\sqrt{2\rho}$. This implies that $|u|_{\infty}\leq c\sqrt{2\rho}$, where $c>0$ is the best constant in the embedding $H^{m}(\mathbb{R})\hookrightarrow L^{\infty}(\mathbb{R})$. Let $\varepsilon>0$: from (F0) and (F1) there exists $C=C(\varepsilon,\rho)>0$ such that $H(t)\leq(\eta+\varepsilon)t^{2+4m}+Ct^{4+4m}$ for every $t\in\overline{B(0,c\sqrt{2\rho})}$. Consequently, from (1.6) and the fact that $u\in\mathcal{M}$,
$2m|u^{(m)}|_{2}^{2}$ $=\int_{\mathbb{R}}H(u)\,\mathrm{d}x\leq(\eta+\varepsilon)|u|_{2+4m}^{2+4m}+C|u|_{4+4m}^{4+4m}$
$\leq(\eta+\varepsilon)C_{2+4m}^{2+4m}|u^{(m)}|_{2}^{2}|u|_{2}^{4m}+CC_{4+4m}^{4+4m}|u^{(m)}|_{2}^{2+1/m}|u|_{2}^{2+4m-1/m}$
$\leq(\eta+\varepsilon)C_{2+4m}^{2+4m}\rho^{2m}|u^{(m)}|_{2}^{2}+CC_{4+4m}^{4+4m}\rho^{1+2m-1/(2m)}|u^{(m)}|_{2}^{2+1/m},$
and we conclude taking $\varepsilon>0$ sufficiently small in view of (1.8). ∎
Lemma 3.6.
If (F0)–(F4) and (1.8) hold, then $\inf_{\mathcal{D}\cap\mathcal{M}}J>0$.
Proof.
We begin by showing the existence of $\delta>0$ such that
$\left(2m-\eta C_{2+4m}^{2+4m}\rho^{2m}\right)|u^{(m)}|_{2}^{2}\leq 8mJ(u)$
for all $u\in\mathcal{D}\cap H^{m}(\mathbb{R})$ with $|u^{(m)}|_{2}\leq\delta$. We can assume $\delta\leq 1$, hence there exists $c_{\rho}>0$ such that $|u|_{\infty}\leq c_{\rho}$ for all $u$ as above. From (F0), (F1), and (F4), for every $\varepsilon>0$ there exists $C=C(\varepsilon,\rho)>0$ such that
$F(t)\leq\left(\frac{\eta}{4m}+\varepsilon\right)t^{2+4m}+Ct^{4+4m}\quad\forall t\in\overline{B(0,c_{\rho})}.$
This and (1.6) yield
$\int_{\mathbb{R}}F(u)\,\mathrm{d}x$ $\leq\left(\frac{\eta}{4m}+\varepsilon\right)|u|_{2+4m}^{2+4m}+C|u|_{4+4m}^{4+4m}$
$\leq\left[\left(\frac{\eta}{4m}+\varepsilon\right)C_{2+4m}^{2+4m}\rho^{2m}+CC_{4+4m}^{4+4m}\rho^{1+2m-1/(2m)}|u^{(m)}|_{2}^{1/m}\right]|u^{(m)}|_{2}^{2}$
$\leq\left[\left(\frac{\eta}{4m}+\varepsilon\right)C_{2+4m}^{2+4m}\rho^{2m}+CC_{4+4m}^{4+4m}\rho^{1+2m-1/(2m)}\delta^{1/m}\right]|u^{(m)}|_{2}^{2},$
which implies, in turn,
$8mJ(u)=4m|u^{(m)}|_{2}^{2}-8m\int_{\mathbb{R}}F(u)\,\mathrm{d}x\\ \geq\left[4m-\left(2\eta+8m\varepsilon\right)C_{2+4m}^{2+4m}\rho^{2m}-8mCC_{4+4m}^{4+4m}\rho^{1+2m-1/(2m)}\delta^{1/m}\right]|u^{(m)}|_{2}^{2}\\ =\left[2\left(2m-\eta C_{2+4m}^{2+4m}\rho^{2m}\right)-8mC_{2+4m}^{2+4m}\rho^{2m}\varepsilon-8mCC_{4+4m}^{4+4m}\rho^{1+2m-1/(2m)}\delta^{1/m}\right]|u^{(m)}|_{2}^{2}.$
The claim then holds true taking
$\varepsilon=\frac{2m-\eta C_{2+4m}^{2+4m}\rho^{2m}}{16mC_{2+4m}^{2+4m}\rho^{2m}}\quad\text{and}\quad\delta=\min\left\{1,\left(\frac{2m-\eta C_{2+4m}^{2+4m}\rho^{2m}}{16mCC_{4+4m}^{4+4m}\rho^{1+2m-1/(2m)}}\right)^{m}\right\}.$
Now let $u\in\mathcal{D}\cap\mathcal{M}$ and set $s=\delta/|u^{(m)}|_{2}$ and $v=s\star u$ so that $v\in\mathcal{D}\cap H^{m}(\mathbb{R})$ and $|v^{\prime}|_{2}=\delta$. Then, Lemma 3.2 yields
$J(u)\geq J(v)\geq\frac{2m-\eta C_{2+4m}^{2+4m}\rho^{2m}}{8m}\delta^{2}.\qed$
Lemma 3.7.
If (F0)–(F4) and (1.8) hold, then $J$ is coercive over $\mathcal{D}\cap\mathcal{M}$.
Proof.
We refer to the proofs of [4, Lemma 2.4] or [11, Lemma 2.5 (iv)], which are similar. ∎
We need the following version of Lions’s Lemma (cf. [19, Lemma 3.1]):
Lemma 3.8.
Let $G\in\mathcal{C}(\mathbb{R})$ such that $G(s)=o(s^{2})$ as $u\to 0$. If $(u_{n})\subset H^{m}(\mathbb{R})$ is bounded and satisfies
$\lim_{n}\max_{y\in\mathbb{R}}\int_{y-r}^{y+r}u_{n}^{2}\,\mathrm{d}x=0$
for some $r>0$, then $\lim_{n\to+\infty}\int_{\mathbb{R}}|G(u_{n})|\,\mathrm{d}x=0$.
Proof.
Let $M>0$ be such that $\sup_{n}|u_{n}|_{\infty}<M$, and take any $\varepsilon>0$ and $p>2$. Then, we find $0<\delta<M$ and $c_{\varepsilon}>0$ such that
$\begin{cases}|G(s)|\leq\varepsilon s^{2}&\text{ if }|s|\in[0,\delta],\\ |G(s)|\leq c_{\varepsilon}|s|^{p}&\text{ if }|s|\in(\delta,M].\end{cases}$
Hence, in view of Lions’ lemma [15, Lemma I.1], we get
$\limsup_{n\to\infty}\int_{\mathbb{R}^{2}}|G(u_{n})|\,\mathrm{d}x\leq\varepsilon\limsup_{n\to\infty}\int_{\mathbb{R}^{2}}u_{n}^{2}\,\mathrm{d}x.$
Letting $\varepsilon\to 0^{+}$, we conclude. ∎
Using Lemma 3.8 and proceeding as in [19, Proof of Theorem 1.4], we have the following result in the spirit of [9].
Proposition 3.9.
Suppose that $(u_{n})\subset H^{m}(\mathbb{R})$ is bounded. Then, there are sequences $(\widetilde{u}_{i})_{i=0}^{\infty}\subset H^{m}(\mathbb{R})$, $(y_{n}^{i})_{i=0}^{\infty}\subset\mathbb{R}$ for every $n$, such that $y_{n}^{0}=0$, $|y_{n}^{i}-y_{n}^{j}|\to\infty$ as $n\to\infty$ for $i\neq j$, and passing to a subsequence, the following conditions hold for every $i\geq 0$:
$u_{n}(\cdot+y_{n}^{i})\rightharpoonup\widetilde{u}_{i}\text{ in }H^{m}(\mathbb{R})\text{ as }n\to\infty,$
(3.4) $\lim_{n\to\infty}\int_{\mathbb{R}}|u_{n}^{(m)}|^{2}\,\mathrm{d}x=\sum_{j=0}^{i}\int_{\mathbb{R}}|\widetilde{u}_{j}^{(m)}|^{2}\,\mathrm{d}x+\lim_{n\to\infty}\int_{\mathbb{R}}|(v^{i}_{n})^{(m)}|^{2}\,\mathrm{d}x,$
where $v_{n}^{i}:=u_{n}-\sum_{j=0}^{i}\widetilde{u}^{j}(\cdot-y_{n}^{j})$, and
(3.5) $\limsup_{n\to\infty}\int_{\mathbb{R}}G(u_{n})\,\mathrm{d}x=\sum_{j=0}^{\infty}\int_{\mathbb{R}}G(\widetilde{u}_{j})\,\mathrm{d}x$
for any function $G\colon\mathbb{R}\to[0,\infty)$ of class $\mathcal{C}^{1}$ such that $|G^{\prime}(s)|=\mathcal{O}(|s|)$ as $s\to 0$ and $G(s)=o(s^{2})$ as $s\to 0$.
Lemma 3.10.
If (F0)–(F4) and (1.8) hold, then $\inf_{\mathcal{D}\cap\mathcal{M}}J$ is attained.
Proof.
Let $(u_{n})_{n}\subset\mathcal{D}\cap\mathcal{M}$ such that $\lim_{n}J(u_{n})=\inf_{\mathcal{D}\cap\mathcal{M}}J$. From Lemma 3.7, we know that $u_{n}$ is bounded in $H^{m}(\mathbb{R})$. Moreover, Proposition 3.9 applied to the function $H$ implies the existence of the sequences $(\widetilde{u}_{i})_{i=0}^{\infty}\subset H^{m}(\mathbb{R})$ satisfying (3.4) and (3.5). Our first claim is that
(3.6) $\exists i\in\mathbb{N}\text{ such that }0<\int_{\mathbb{R}}|\widetilde{u}_{i}^{(m)}|^{2}\,\mathrm{d}x\leq\frac{1}{2m}\int_{\mathbb{R}}H(\widetilde{u}_{i})\,\mathrm{d}x.$
Let $I:=\{i\in\mathbb{N}:\widetilde{u}_{i}\neq 0\}$. In view of Lemma 3.5, (3.5), and the fact that $u_{n}\in\mathcal{M}$, it is clear that $I\neq\emptyset$. Arguing by contradiction, assume that
$\int_{\mathbb{R}}|\widetilde{u}_{i}^{(m)}|^{2}\,\mathrm{d}x>\frac{1}{2m}\int_{\mathbb{R}}H(\widetilde{u}_{i})\,\mathrm{d}x\quad\forall i\in I.$
Then, (3.4) and (3.5) imply that
$\limsup_{n\to\infty}\frac{1}{2m}\int_{\mathbb{R}}H(u_{n})\,\mathrm{d}x$ $=\limsup_{n\to\infty}\int_{\mathbb{R}}|u_{n}^{(m)}|^{2}\,\mathrm{d}x\geq\sum_{j=0}^{\infty}\int_{\mathbb{R}}|(\widetilde{u}_{j})^{(m)}|^{2}\,\mathrm{d}x$
$=\sum_{j\in I}\int_{\mathbb{R}}|(\widetilde{u}_{j})^{(m)}|^{2}\,\mathrm{d}x>\sum_{j\in I}\frac{1}{2m}\int_{\mathbb{R}}H(\widetilde{u}_{j})\,\mathrm{d}x$
$=\limsup_{n\to\infty}\frac{1}{2m}\int_{\mathbb{R}}H(u_{n})\,\mathrm{d}x,$
which is impossible. Consequently, (3.6) holds for some $i\in I$, and note that $r_{i}\geq 1$, where
$r_{i}^{2}=\frac{\int_{\mathbb{R}}H(\widetilde{u}_{i})\,\mathrm{d}x}{2m|\widetilde{u}_{i}^{(m)}|_{2}^{2}}.$
Additionally, a straightforward computation shows that $u:=\widetilde{u}_{i}(r_{i}\cdot)\in\mathcal{M}$. In fact, $u\in\mathcal{D}\cap\mathcal{M}$ because $r_{i}\geq 1$, therefore, from (F4) and Fatou’s Lemma, we have
$\inf_{\mathcal{D}\cap\mathcal{M}}J$ $\leq J(u)=\int_{\mathbb{R}}\frac{1}{4m}H(u)-F(u)\,\mathrm{d}x=\frac{1}{r_{i}}\int_{\mathbb{R}}\frac{1}{4m}H(\widetilde{u}_{i})-F(\widetilde{u}_{i})\,\mathrm{d}x$
$\leq\int_{\mathbb{R}}\frac{1}{4m}H(\widetilde{u}_{i})-F(\widetilde{u}_{i})\,\mathrm{d}x$
$\leq\liminf_{n\to+\infty}\int_{\mathbb{R}}\frac{1}{4m}H\bigl{(}u_{n}(x+y_{n}^{i})\bigr{)}-F\bigl{(}u_{n}(x+y_{n}^{i})\bigr{)}\,\mathrm{d}x$
$\leq\lim_{n\to+\infty}\int_{\mathbb{R}}\frac{1}{4}H(u_{n})-F(u_{n})\,\mathrm{d}x=\lim_{n\to+\infty}J(u_{n})=\inf_{\mathcal{D}\cap\mathcal{M}}J,$
which implies that $r_{i}=1$ and $u=\widetilde{u}_{i}$ minimises $J$ over $\mathcal{D}\cap\mathcal{M}$. ∎
Proof of Theorem 1.5.
The first part follows from Lemma 3.10. Now, let $u$ be the minimiser of $J$ over $\mathcal{D}\cap\mathcal{M}$ given therein. Our first claim is that the functional $\bigl{(}\Phi^{\prime}(v),M^{\prime}(v)\bigr{)}\colon H^{m}(\mathbb{R})\to\mathbb{R}^{2}$ is surjective for every $v\in\mathcal{S}\cap\mathcal{M}$, where $\Phi(v):=|v|_{2}^{2}$ and
$M(v):=\int_{\mathbb{R}}\Bigl{(}|v^{(m)}|^{2}-\frac{1}{2m}H(v)\Bigr{)}\,\mathrm{d}x.$
Indeed, given $v\in\mathcal{S}\cap\mathcal{M}$, we consider the curve $(0,\infty)\ni s\mapsto s\star v\in\mathcal{S}$, and the function $\psi_{v}(s)=M(s\star v)=s\varphi^{\prime}_{v}(s)$. We notice that the curve $s\mapsto s\star v$ is not tangent at $v$ to the manifold $\mathcal{M}$, since otherwise we would obtain $\psi^{\prime}_{v}(1)=\varphi^{\prime\prime}_{v}(1)=0$ in contradiction with Remark 3.4). Thus, the manifolds $\mathcal{S}$ and $\mathcal{M}$ do not have the same tangent plane at $v$. Hence, from [20, Proposition A.1], there exist $\lambda\geq 0$ and $\theta\in\mathbb{R}$ such that
(3.7) $(1+\theta)\Bigl{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}\Bigr{)}^{m}u+\lambda u=F^{\prime}(u)+\frac{\theta}{4m}H^{\prime}(u).$
If $\theta=-1$, then (F3) and (F4), together with (3.7), imply
$0\leq\lambda\int_{\mathbb{R}}|u|^{2}\,\mathrm{d}x$ $=\int_{\mathbb{R}}\biggl{(}F^{\prime}(u)u-\frac{1}{4m}H^{\prime}(u)u\biggr{)}\,\mathrm{d}x<\int_{\mathbb{R}}\biggl{(}F^{\prime}(u)u-\Bigl{(}\frac{1}{2m}+1\Bigr{)}H(u)\biggr{)}\,\mathrm{d}x$
$=\int_{\mathbb{R}}\biggl{(}2F(u)-\frac{1}{2m}H(u)\biggr{)}\,\mathrm{d}x\leq 0,$
a contradiction. As a consequence, from Proposition 3.1, $u$ satisfies also the Pohožaev and Nehari identity associated with (3.7), whence
$(1+\theta)\int_{\mathbb{R}}|u^{(m)}|^{2}\,\mathrm{d}x=\frac{1}{2m}\int_{\mathbb{R}}H(u)+\frac{\theta}{4m}\bigl{(}H^{\prime}(u)u-2H(u)\bigr{)}\,\mathrm{d}x.$
Since $u\in\mathcal{M}$, we obtain
$\theta\int_{\mathbb{R}}\bigl{(}H^{\prime}(u)u-(2+4m)H(u)\bigr{)}\,\mathrm{d}x=0,$
whence $\theta=0$ in view of (F3) and (1.7). This proves that $\bigl{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}\bigr{)}^{m}u+\lambda u=F^{\prime}(u)$. Since $\lambda=0$ if $u\in\mathcal{D}\setminus\mathcal{S}$, we only need to verify that $\lambda>0$. Indeed, if by contradiction, $\lambda=0$, then the Pohožaev identity yields
$\int_{\mathbb{R}}F(u)\,\mathrm{d}x=-\frac{1}{2}\int_{\mathbb{R}}|u^{(m)}|^{2}\,\mathrm{d}x<0,$
which contradicts (F4). ∎
Acknowledgements
This work was supported by the Thematic Research Programme “Variational and geometrical methods in partial differential equations”, University of Warsaw, Excellence Initiative - Research University. J.S. is a member of GNAMPA (INdAM) and was supported by the programme Borse di studio per l’estero dell’Istituto Nazionale di Alta Matematica and the GNAMPA project Metodi variazionali e topologici per alcune equazione di Schrödinger nonlineari. The research project of Panayotis Smyrnelis is implemented in the framework of H.F.R.I call “Basic research Financing (Horizontal support of all Sciences)” under the National Recovery and Resilience Plan “Greece 2.0” funded by the European Union - NextGenerationEU (H.F.R.I. Project Number: 016097).